40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6913     Entry time: Wed Jul 4 20:13:46 2012
Author: yuta 
Type: Bureaucracy 
Category: Locking 
Subject: PRC commissioning plan 

Issues in PRC:
  1. Power recycling gain is too low (~ 15 instead of 40, according to Kiwamu).
  2. Mode matching to both arms are ~90%(see #6859), but PRC has terrible mode.
       Clipping/flipping in PRC?
  3. From cameras, beam spot moves so much when PRMI is locked.
       Alignment? Mirrors(especially PR2/3) moves too much?
  4. Error signals are glitchy when PRMI is locked.
       Servo design? Mirrors moves too much?

What we have learned from the vent:

  1. PRM, PR2, PR3 was not flipped.
  2. Their suspensions looked OK, too.
  3. We noticed clipping at BS and Faraday. They must be avoided when tip-tilts are installed on next vent.

  4. Took useful photos for next vent. Positions of green optics on optical layout CAD must be updated.
  5. It is not so difficult to recover the IFO state after cycling the vacuum if we use attenuator setup using PBS (see elog #6892).  But, of course, we need plans before cycling.

Commissioning Plan:
  - measure PRMI power recycling gain from POP
  - FPMI using ALS
  - measure PRFPMI power recycling gain from TRY/X
  - correlation between beam spot motion at POP camera and glitch
  - correlation between PR2/PR3 motion and glitch (how can we measure PR2/3 motion? set up oplevs?)
  - mode scan for PRC, using AS AUX laser
  - beam profile measurement at REFL,POP
  - refine servo design of MICH and PRCL

ELOG V3.1.3-